# The Fall of an Algorithm: Characterizing the Dynamics Toward Abandonment

### v1

@@ 4. The Dynamics Toward Algorithm Abandonment
We explore the dynamics of abandonment by proposing a process-oriented view of the events leading to abandonment. We visualized these dynamics for individual cases by creating a timeline of events (shown in Figure 1) that occurred until a system was abandoned (or until now for systems that are in use). In our thematic analysis, we found that most algorithms in our case set were abandoned after they were deployed. However, our case set also includes cases of algorithms abandoned earlier in their development lifecycle. These cases provide meaningful comparison points to consider how algorithmic harms might be prevented.
The 6 D’s of abandonment. In our analysis of event timelines for the 40 identified cases, the process leading to abandonment involved events that fell into six iterative phases of (1) discovery, which initiated critique and calls to abandon the system by the public, impacted communities, or their advocates, (2) diagnosis, where these actors gathered further evidence about the system and its impacts, (3) dissemination, where these actors amplified their message to a wider audience, (4) dialogue between critics of the system and algorithm owners, until they reached a (5) decision about whether to repair or abandon the system. We discuss how many abandoned systems still continued to impact peoples’ lives in the final (6) death phase, in which algorithm owners transitioned away from the algorithm and those impacted sought justice for past harms. Note that all six phases may occur in any stage of an algorithm’s lifecycle (e.g., design, development, usage), and are still relevant for algorithms that are not yet developed or deployed. In practice, the “fall” of an algorithm often proceeds iteratively and non-linearly throughout these six phases. Similarly, not all six phases may occur for every case. Regardless, this process-oriented view provides a useful lens to examine what types of events were influential across a range of cases.
@@ 4.1. Discovery
Many cases had a catalyzing event where information that was previously unknown about the algorithm was made public and initiated wide-scale critique. We group these discoveries into two categories: (1) discovering the existence of a previously obscure algorithmic system; and (2) discovering potential or realized algorithmic harms.
@@
Discovery of existence. In some cases, criticism began once the algorithm’s mere existence was made public. While algorithm owners could in theory provide public notice of their intentions to develop or use an algorithm in early stages of its lifecycle, many did not provide notice until the system was already in use. For example, the City of Pittsburgh “experimented with” using a predictive policing algorithm for over a year before the public was made aware of its existence through the release of a public report (in Pittsburgh, 2020a). Several other algorithms such as the New York Police Department’s contract with facial recognition software vendor, Clearview AI (Haskins, 2021), evaded critique for months due to a lack of public notice. Discoveries of an algorithms’ existence sparked criticism about a lack of community engagement and oversight (in Pittsburgh, 2020a) and the values encoded in algorithms (Hill, 2020a; log, 2020; Khan and White, 2021).
@@
Discovery of algorithmic harm. In other cases, initial critiques began after a discovery of potential or realized algorithmic harms. To understand what types of harms motivated calls to abandonment, we mapped all mentions of harms in case materials to the five major types of sociotechnical harms in Shelby et al. (2023) taxonomy (defined in Appendix D), i.e., representational, allocative, quality-of-service, interpersonal, and social system harms. Overall, the algorithms in our case set spanned all five of Shelby et al. (2023)’s harm categories. Algorithms deployed in similar domains resulted in similar harms, e.g., chatbots could demean and stereotype (Kwon and Yun, 2021; Kraft, 2016), and resource allocation algorithms could result in opportunity loss (Ryan-Mosley, 2023a; Hub, 2022a). Several different types of harm often co-occurred for each single algorithmic system.444For example, San Diego’s TACIDS algorithm (System, 2021; Williams, 2015) matched photographs of suspects against mugshot databases, where people with previous law enforcement encounters were more likely to be (falsely) flagged (representational harm). TACIDS was used to aid immigration enforcement efforts that separated families (interpersonal harm), and had implications for residents’ constitutional rights (social system harms). Finally, critics identified harms that implicated more than just the algorithms’s functionality, or ability to work as intended (Raji et al., 2022). In other words, simply improving the algorithm to be more accurate or to reduce what Shelby et al. (2023) deem “Quality-of-Service” performance disparities would not redress all harms. A more effective surveillance technology, for example, could still contribute to civil rights violations or interpersonal harms, such as accelerating incarceration, or family separation at scale (Wang et al., 2022; System, 2021). While many cases we considered did involve dysfunctional algorithms, the remaining taxonomy categories remind us that other harms exist too.
@@ 4.2. Diagnosis
Examining the output: algorithm auditing. Critics attempted to better understand the behavior of an algorithm by interacting with or probing it directly to observe its output. Many such audits involved querying the algorithm on several (real or simulated) inputs, and measuring properties of its output. Large-scale targeted audits were a highly effective tool that precipitated decisions to abandon an algorithm. For example, many organizations referenced the Gender Shades audit (which measured racial and gender accuracy disparities of commercial facial recognition algorithms (Buolamwini and Gebru, 2018)) when they later instituted moratorium policies on the sales or usage of facial recognition technology (Buolamwini and Gebru, 2023). While large-scale audits were highly effective, in several cases, critics faced significant obstacles to conduct such audits. One dimension that significantly influenced the auditability of an algorithm was access to the model, as algorithms that were administered by a decision-maker internal to the organization (e.g., a social worker assessing child welfare cases (Ho and Burke, 2022)) could not be queried. Independent auditors also often lacked access to realistic and representative data of the population of interest. In some cases, the public was even unaware of what information the algorithm took as input to make decisions. For example, in Jordan, government officials refuse to release the “socio-economic indicators” of poverty that an algorithm which allocated cash assistance to citizens took into consideration (Watch, 2023).
@@
To overcome these obstacles, in some cases critics worked together to collectively develop work-around strategies. One strategy to obtain access to realistic input data was to design coordinated data collection efforts from consenting people who would be or were already impacted by an algorithmic system. In Missouri, a coalition of legal aid organizations and third-party researchers partnered with home care nurses to fill out assessments to estimate a proposed home care allocation algorithms’ impact for 1200 patients (Weil and Edwards, 2023). In Spain, an independent nonprofit conducted phone interviews to understand the experiences of women who completed an algorithmic gender-based violence risk assessment that determined the police and legal resources allocated to them (eti, 2022). Successful data donation initiatives gave auditors access to population-level data that was critical to understanding the algorithm’s behavior.
@@
Examining the process: design decisions. Algorithmic systems do not just “come about”, but are rather the product of a series of decisions made by designers, developers, and deployers (Selbst et al., 2024; Metcalf et al., 2023). Critics attempted to understand algorithmic systems by examining the important decisions that brought them into being. For example, concerned Carnegie Mellon University students who opposed Pittsburgh’s use of a predictive policing model filed a Right to Know request that revealed that the algorithm was rolled out without any meaningful consultation with the communities it was used to police (in Pittsburgh, 2020a). Several advocacy groups have created resources to help people collect information about human decisions and processes surrounding algorithmic systems. For algorithms used by public agencies, organizations including the Electronic Frontiers Foundation (“EFF”) (eff, 2023) and Lucy Parsons Labs (lpl, 2023) have developed resources to help people navigate filing public records requests (which in some cases only resulted in partial information being released (in Pittsburgh, 2020a; Marx and Irani, 2021)).
@@ 4.3. Dissemination
Broader dissemination was a common strategy employed by critics to raise public awareness and put public pressure on algorithm owners. Dissemination sometimes caught the attention of elected officials or regulators who could apply pressure or exercise authority from above (Ho and Burke, 2022; Warren, 2023; Ho and Burke, 2023). Some campaigns led by advocates of abandonment took place in-person (offline), whereas in other cases campaigns took place online. Other cases were supported by global media campaigns, such as the #BanTheScan coalition supporting grassroots activist groups’ efforts to ban government use of facial recognition technologies in NYC, Jerusalem, and Hyderabad (amn, 2022). In many instances (Liban, 2018; in Pittsburgh, 2020b; ap2, 2022; was, [n. d.]), only after the visibility of criticism increased did algorithm owners respond to and acknowledge the critique.
Online vs. Offline. While less common, a small number of cases involved in-person public demonstrations. All of the public demonstrations in our case set were organized to oppose systems that were intended to be deployed in a specific locality (e.g., in a particular city (Chang, 2018)). Many such demonstrations were organized and attended by people who were directly impacted by the algorithm, such as student-organized protests against algorithmically assigned exam grades in the UK (Porter, 2020), or a resident-led protest of a proposed facial recognition algorithm held on the sidewalk outside of their apartment (Elizalde and Gartland, 2019). In contrast, criticism of tools with spatially diffuse user bases took place online. Several cases gained visibility through high engagement on online platforms, such as Twitter, Instagram, or YouTube by going “viral”, such as users’ interactions with NEDA’s Tessa Bot (Wells, 2023), or the Korean chatbot Lee-Luda (Kwon and Yun, 2021). We found that highly visible cases (as approximated through engagement metrics, such as high reaction or re-posts counts) were significantly more likely to receive a timely response from the algorithm owner. Social media campaigns involved broad participation from a wide range of stakeholders, who shared reactions to other posts, shared their own interactions with the system, or simply liked or shared to boost its visibility.
Getting press. In several cases, media coverage of criticism played a critical role in the fall of an algorithm. As one example, Oregon’s DHS decided just two weeks after an Associated Press story (Ho and Burke, 2022) to halt a pilot on child maltreatment risk prediction that had been in progress for four years. US Senator Ron Wyden had “reached out” to the department immediately after the story to “ask questions about [the tool’s] racial bias” (ap2, 2022). Similar reporting efforts helped activist efforts gain momentum, and caught the attention of those in positions of power who applied pressure from above.
@@ 4.4. Dialogue
When criticism about an algorithm became more publicly visible, actors who held power over the algorithm were called to respond in a public dialogue with their critics. These dialogues occurred in sites ranging from those that are more formal or compulsory for algorithm owners (e.g., in court (Hub, 2022a, b; van Veen, 2020; Pierson and Pierson, 2023; Mari, 2022)), to those that were more informal (e.g., on social media (Yee et al., 2021; Dickey, 2020; Vincent, [n. d.]; Wells, 2023; https://www.theguardian.com/profile/justinmccurry, [n. d.])). Some algorithm owners took weeks, months, or even years to acknowledge critics and respond at all. In this section, we examine how algorithm owners (dis)engaged with criticism about the algorithm. While some algorithms were primarily criticized by the public, other algorithms found support from other actors beyond their owners, and sparked a broader discourse about whether or not they should exist.
@@
Denial vs. acceptance of algorithmic harm. When presented with evidence of algorithmic harm, some algorithm owners responded by discrediting the evidence presented to them, either by creating counter-evidence (Eubanks, 2018) or ad-hominem attempts to discredit the critic (Hao, 2020; Xiang, 2023). For example, when an Instagram user shared an infographic that described a harmful and triggering interaction with NEDA’s Tessa Wellness chatbot555https://www.instagram.com/p/Cs1jp1pPkOs/, NEDA’s VP of Communications accused her of lying in the post’s comment section (Xiang, 2023). Interestingly, in many cases (Hao, 2020; Xiang, 2023), initial denials were later retracted or withdrawn due to public scrutiny. After the NEDA VP’s comment sparked outrage from other users, she deleted her comment and posted that she “retracted her statement” (Xiang, 2023). This pattern of denial followed by retraction is a broader trend in responses to attempts for algorithmic accountability (Raji and Buolamwini, 2022). In contrast, rather than resist, some algorithm owners immediately acknowledged the validity of the critiques. For example, when Twitter users began to investigate if its image cropping algorithm was racially biased (Yee et al., 2021), Twitter immediately issued a public statement that they were ”prioritizing work” to remedy bias in their algorithms (Agrawal and Davis, 2020). Cooperative responses that immediately acknowledged critiques allowed organizations to take actions (including abandonment) to redress harms more quickly.
@@
The public vs. private dialogue. One dimension that influenced whether an algorithm was abandoned was whether algorithm owners responded in private. Cases that were resolved interpersonally rather than in public were significantly less likely to result in critiques being resolved or the system being abandoned. For example, while Zoom users observed in their personal experiences (Shen et al., 2021) that Zoom’s virtual background feature tended to be more dysfunctional for users with darker skin, prompting some to tweet about it, Zoom did not respond until one of these tweets666https://twitter.com/colinmadland/status/1307111816250748933 went viral. Zoom Corporate replied to the Tweet to invite its author to a call. Later that week, the author of the tweet wrote that he had “had a great convo with [Zoom’s Chief DEI Officer] […] I have no reason to doubt that they are serious about fixing problems with virtual backgrounds […] the problem is deep and nuanced”. Yet, such private resolutions still leave the broader public in the dark about what commitments the algorithm owner has made.
@@ 4.5. Decision
Those who held power over the algorithm could make decisions to intervene on its design, development, or usage. We group these decisions about the algorithm into 4 categories based on their outcome: whether the algorithm was not abandoned (in which the algorithm owners could (1) repair the algorithm, or (2) make no changes), or was abandoned (3) by the algorithm owner’s choice, or (4) by law or policy. In some cases, multiple decisions were made, e.g., many algorithm owners first attempted to repair an algorithm before later deciding to abandon it.
@@
Repair. Algorithm owners took several actions to repair or remediate potential harms, including fixing or updating dysfunctional technology (Liban, 2018; Vaithianathan et al., 2019; Ross, 2021) or creating better infrastructure or policies around its use (Dastin, 2020; Sisitzky and Schaefer, 2021; Ross, 2021). For example, when independent investigations revealed flaws in Epic’s sepsis prediction model (which was deployed in several hundred U.S. hospitals) (Wong et al., 2021), the company completely “overhauled” its original algorithm and released a second version that used completely different input and target variables, and recommended that the model be trained on a hospital’s own data rather than being applied off-the-shelf (Ross, 2021). But repairs that improved the technology by making it more functional or less biased did not redress harms that resulted from the algorithms’ existence.
@@
Abandoned by the algorithm owner’s choice. In our analysis, the majority of systems that were abandoned were abandoned by choice, i.e., by a decision made by an organization that developed or used the algorithm. To name a few, Korean start-up ScatterLab disabled their Lee-Luda chatbot (Kwon and Yun, 2021), Twitter decided to stop using an algorithm to crop users’ images (Chowdhury, 2021), and leadership at Oregon’s Department of Human Services decided to stop using a child welfare call screening algorithm (ap2, 2022). Relatively few algorithm owners released public statements with explanations or named reasons why they decided to abandon the algorithm. Public statements that we found referenced algorithmic harms as a primary reason why the system was abandoned (Lee, 2016; Chowdhury, 2021; Kwon and Yun, 2021; Hoover, 2023).
@@
Abandoned through policy. Algorithms overseen by uncooperative algorithm owners could still be abandoned through another governing body creating or enforcing applicable laws. Several algorithms in our case set were deemed illegal through the interpretation of existing laws (including laws written without an algorithm in mind) in a court of law. For example, a Dutch court ruled the government to stop using an algorithm that falsely predicted that thousands of low-income residents had committed benefits fraud, after finding that the algorithm violated citizens’ right to privacy in the European Convention on Human Rights (aw2, 2020). U.S. courts ruled MedicAid benefits allocation algorithms in violation of existing laws (i.e., due process law and the Administrative Procedure Act) in both Arkansas and Idaho (Calo and Citron, 2021; Eppink, 2023). Other algorithms were banned by existing governing bodies founded to protect citizens’ rights, such as the U.S. Federal Trade Commission issuing a five-year ban on Rite Aid’s use of facial recognition because of the technology’s risks to consumers (ftc, 2023). However, all past cases of systems abandoned through executive or judicial enforcement took a significantly longer time (e.g., several years (Liban, 2018; Eppink, 2023)) before the decision to abandon the algorithm was reached. These cases also required significant investment - e.g., the legal team representing Idaho plaintiffs spent over $40,000 on expert consultations and 2,000 attorney hours to secure a settled agreement (Eppink, 2023).

### v2

@@ 4. The Dynamics Toward Algorithm Abandonment
We explore the dynamics of abandonment by proposing a process-oriented view of the events leading to abandonment. We visualized these dynamics for individual cases by creating a timeline of events (shown in Figure 1) that occurred until a system was abandoned (or until the time of writing for systems that were still in use). In our thematic analysis, we found that most algorithms in our case set were abandoned after they were deployed. However, our case set also includes cases of algorithms abandoned earlier in their development lifecycle. These cases provide meaningful comparison points to consider how algorithmic harms might be anticipated and prevented before a system is released.
The 6 D’s of abandonment. In our analysis of event timelines for the 40 identified cases, the process leading to abandonment involved events that fell into six iterative phases of (1) discovery, which initiated critique and calls to abandon the system by the public, impacted communities, or their advocates, (2) diagnosis, where these actors gathered further evidence about the system and its impacts, (3) dissemination, where these actors amplified their concerns to a wider audience, (4) dialogue between critics of the system and algorithm owners, until they reached a (5) decision about whether to repair or abandon the system. We discuss how many abandoned systems still continued to impact peoples’ lives in the final (6) death phase, in which algorithm owners transitioned away from the algorithm and those impacted sought justice for past harms. Note that all six phases may occur in any stage of an algorithm’s lifecycle (e.g., design, development, usage), and are still relevant for algorithms that are not yet developed or deployed. In practice, the “fall” of an algorithm often proceeds iteratively and non-linearly throughout these six phases. Similarly, not all six phases may occur for every case. Regardless, this process-oriented view provides a useful lens to examine what types of events were influential across a range of cases.
@@ 4.1. Discovery
Many cases had a catalyzing event where information that was previously unknown about the algorithm was made public and initiated wide-scale critique. We group these discoveries into two categories: (1) discovering the existence of a previously obscured algorithmic system; and (2) discovering potential or realized algorithmic harms.
@@
Discovery of existence. In some cases, criticism began once the algorithm’s mere existence was made public. While algorithm owners could in theory provide public notice of their intentions to develop or use an algorithm in early stages of its lifecycle, many did not provide notice until the system was already in use. For example, the City of Pittsburgh “experimented with” using a predictive policing algorithm for over a year before the public was made aware of its existence through the release of a public report (in Pittsburgh, 2020a). Several other algorithms such as the New York Police Department’s contract with facial recognition software vendor Clearview AI (Haskins, 2021) evaded critique for months due to a lack of public notice. Discoveries of an algorithms’ existence often sparked criticism about a lack of community engagement and oversight (in Pittsburgh, 2020a) and the values they encoded (e.g., through their task formulation or deployers’ stated goals) (Hill, 2020a; log, 2020; Khan and White, 2021).
@@
Discovery of algorithmic harm. In other cases, initial critiques began after a discovery of potential or realized algorithmic harms. To understand what types of harms motivated calls to abandonment, we mapped all mentions of harms in case materials to the five major types of sociotechnical harms in Shelby et al. (2023): representational, allocative, quality-of-service, interpersonal, and social system harms. We provide complete definitions and examples of each harm type Appendix Table 2. In summary, the algorithms in our case set spanned all five of Shelby et al. (2023)’s harm categories. Algorithms deployed in similar domains resulted in similar harms, e.g., chatbots could demean and stereotype (Kwon and Yun, 2021; Kraft, 2016), and resource allocation algorithms could result in opportunity loss (Ryan-Mosley, 2023a; Hub, 2022a). Several different types of harm often co-occurred for each single algorithmic system.444For example, San Diego’s TACIDS algorithm (System, 2021; Williams, 2015) matched photographs of suspects against mugshot databases, where people with previous law enforcement encounters were more likely to be (falsely) flagged (representational harm). TACIDS was used to aid immigration enforcement efforts that separated families (interpersonal harm), and had implications for residents’ constitutional rights (social system harms).
Importantly, critics identified harms that implicated more than just the algorithm’s functionality, or ability to work as intended (Raji et al., 2022). In other words, simply improving the algorithm to be more accurate, i.e., to reduce what Shelby et al. (2023) deem “Quality-of-Service” performance disparities, would not redress all harms. A more effective surveillance technology, for example, could still contribute to civil rights violations or interpersonal harms, such as accelerating incarceration or family separation at scale (Wang et al., 2022; System, 2021). While many cases we considered did involve dysfunctional algorithms, the remaining taxonomy categories remind us that other harms exist too.
@@ 4.2. Diagnosis
Examining the output: algorithm auditing. Critics attempted to better understand the behavior of an algorithm by interacting with or probing it directly to observe its output. Many such audits involved querying the algorithm on several (real or simulated) inputs, and measuring properties of its output. Large-scale targeted audits were a highly effective tool that precipitated decisions to abandon an algorithm. For example, many organizations referenced the Gender Shades audit (which measured racial and gender accuracy disparities of commercial facial recognition algorithms (Buolamwini and Gebru, 2018)) when they later instituted moratorium policies on the sale or usage of facial recognition technology (Buolamwini and Gebru, 2023).
While large-scale audits were highly effective, in several cases, critics faced significant obstacles to conduct such audits. One dimension that significantly influenced the auditability of an algorithm was access to the model, as algorithms that were administered by a decision-maker internal to the organization (e.g., a social worker assessing child welfare cases (Ho and Burke, 2022)) could not be queried at will. Independent auditors also often lacked access to realistic and representative data from the population of interest. In some cases, the public was even unaware of what information the algorithm took as input to make decisions. For example, in Jordan, government officials refuse to release the “socio-economic indicators” of poverty that an algorithm which allocated cash assistance to citizens took into consideration (Watch, 2023).
@@
To overcome these obstacles, in some cases critics worked together to collectively develop work-around strategies. One strategy to obtain access to realistic input data was to design coordinated data collection efforts from consenting people who would be or were already impacted by an algorithmic system. In Missouri, a coalition of legal aid organizations and third-party researchers partnered with home care nurses to fill out assessments to estimate a proposed home care allocation algorithm’s impact for 1200 patients (Weil and Edwards, 2023). In Spain, an independent nonprofit conducted phone interviews to understand the experiences of women who completed an algorithmic gender-based violence risk assessment that determined the police and legal resources allocated to them (eti, 2022). Successful data donation initiatives gave auditors access to population-level data that was critical to understanding the algorithm’s behavior.
@@
Examining the process: design decisions. Algorithmic systems do not just “come about”, but are rather the product of a series of decisions made by designers, developers, and deployers (Selbst et al., 2024; Metcalf et al., 2023). Critics attempted to understand algorithmic systems by examining the important decisions that brought them into being. For example, concerned Carnegie Mellon University students who opposed Pittsburgh’s use of a predictive policing model filed a Right to Know request that revealed that the algorithm was rolled out without any meaningful consultation with the communities it was used to police (in Pittsburgh, 2020a). Several advocacy groups have created resources to help people collect information about human decisions and processes surrounding algorithmic systems. For algorithms used by public agencies, organizations including the Electronic Frontier Foundation (“EFF”) (eff, 2023) and Lucy Parsons Labs (lpl, 2023) have developed resources to help people navigate filing public records requests, which in some cases still only resulted in partial information being released (in Pittsburgh, 2020a; Marx and Irani, 2021).
@@ 4.3. Dissemination
Broader dissemination was a common strategy employed by critics to raise public awareness and put public pressure on algorithm owners. Dissemination sometimes caught the attention of elected officials or regulators who could apply pressure or exercise authority from above (Ho and Burke, 2022; Warren, 2023; Ho and Burke, 2023). Some campaigns led by advocates of abandonment took place in-person (offline), whereas in other cases campaigns took place online. Other cases were supported by global media campaigns, such as the #BanTheScan coalition supporting grassroots activist groups’ efforts to ban government use of facial recognition technologies in NYC, Jerusalem, and Hyderabad (amn, 2022). In many instances, only after the visibility of criticism increased did algorithm owners respond to and acknowledge the critique.
Online vs. Offline. While less common, a small number of cases involved in-person public demonstrations. All of the public demonstrations in our case set were organized to oppose systems that were intended to be deployed in a specific locality (e.g., in a particular city (Chang, 2018)). Many such demonstrations were organized and attended by people who were directly impacted by the algorithm, such as student-organized protests against algorithmically assigned exam grades in the UK (Porter, 2020), or a resident-led protest of a proposed facial recognition algorithm held on the sidewalk outside of their apartment (Elizalde and Gartland, 2019). In contrast, criticism of tools with spatially diffuse user bases took place online. Several cases gained visibility through high engagement on online platforms, such as Twitter, Instagram, or YouTube by going “viral”, such as users’ interactions with NEDA’s Tessa Bot (Wells, 2023), or the Korean chatbot Lee-Luda (Kwon and Yun, 2021). We found that highly visible cases (as approximated through engagement metrics, such as high reaction or re-posts counts) were significantly more likely to receive a timely response from the algorithm owner.
Getting press. In several cases, media coverage of criticism played a critical role in the fall of an algorithm. As one example, Oregon’s DHS decided just two weeks after an Associated Press story (Ho and Burke, 2022) to halt a pilot on child maltreatment risk prediction that had been in progress for four years. US Senator Ron Wyden “reached out” to the department immediately after the story to “ask questions about [the tool’s] racial bias” (ap2, 2022). Similar reporting efforts helped activist efforts gain momentum, and sometimes caught the attention of those in positions of power who applied pressure from above.
@@ 4.4. Dialogue
When criticism about an algorithm became more publicly visible, actors who held power over the algorithm were called to respond in a public dialogue with their critics. These dialogues occurred in sites ranging from those that are more formal or compulsory for algorithm owners (e.g., in court (Hub, 2022a, b; van Veen, 2020; Pierson and Pierson, 2023; Mari, 2022)), to those that were more informal (e.g., on social media (Yee et al., 2021; Dickey, 2020; Vincent, [n. d.]; Wells, 2023; https://www.theguardian.com/profile/justinmccurry, [n. d.])). Some algorithm owners took weeks, months, or even years to acknowledge their critics and respond at all. In this section, we examine how algorithm owners (dis)engaged with criticism about the algorithm. While some algorithms were primarily criticized by the public, other algorithms found support from other actors beyond their owners, and sparked a broader discourse about whether or not they should exist.
@@
Denial vs. acceptance of algorithmic harm. When presented with evidence of algorithmic harm, some algorithm owners responded by discrediting the evidence presented to them, either by creating counter-evidence (Eubanks, 2018) or ad-hominem attempts to discredit the critic (Hao, 2020; Xiang, 2023). For example, when an Instagram user shared an infographic that described a harmful and triggering interaction with NEDA’s Tessa Wellness chatbot555https://www.instagram.com/p/Cs1jp1pPkOs/, NEDA’s VP of Communications accused her of lying in the post’s comment section (Xiang, 2023). Interestingly, in many cases (Hao, 2020; Xiang, 2023), initial denials were later retracted or withdrawn due to public scrutiny. After the NEDA VP’s comment sparked outrage from other users, she deleted her comment and posted that she “retracted her statement” (Xiang, 2023). This pattern of denial followed by retraction is a broader trend in responses to attempts for algorithmic accountability (Raji and Buolamwini, 2022).
In contrast, rather than resist, some algorithm owners immediately acknowledged the validity of the critiques. For example, when Twitter users began to investigate if its image cropping algorithm was racially biased (Yee et al., 2021), Twitter immediately issued a public statement that they were “prioritizing work” to remedy bias in their algorithms (Agrawal and Davis, 2020). Cooperative responses that immediately acknowledged critiques allowed organizations to take actions (including abandonment) to redress harms more quickly.
@@
The public vs. private dialogue. One dimension that influenced whether an algorithm was abandoned was whether algorithm owners responded publicly. Cases that were resolved interpersonally rather than in public were significantly less likely to result in critiques being resolved or the system being abandoned. For example, Zoom users observed in their personal experiences (Shen et al., 2021) that Zoom’s virtual background feature tended to be more dysfunctional for users with darker skin, prompting some to tweet about it, but Zoom did not respond until one of these tweets666https://twitter.com/colinmadland/status/1307111816250748933 went viral. Zoom corporate replied to the tweet to invite its author to a call. Later that week, the author of the tweet posted that he had “had a great convo with [Zoom’s Chief DEI Officer] […] I have no reason to doubt that they are serious about fixing problems with virtual backgrounds […] the problem is deep and nuanced”. Yet, such private resolutions still leave the broader public in the dark about what commitments the algorithm owner has made.
@@ 4.5. Decision
Those who held power over the algorithm could make decisions to intervene on its design, development, or usage. We group these decisions about the algorithm into 4 categories based on their outcome: whether the algorithm was not abandoned, in which the algorithm owners could (1) repair the algorithm, or (2) make no changes; or the algorithm was abandoned (3) by the algorithm owner’s choice, or (4) by law or policy. In some cases, multiple decisions were made, e.g., many algorithm owners first attempted to repair an algorithm before later deciding to abandon it.
@@
Repair. Algorithm owners took several actions to repair or remediate potential harms, including fixing or updating dysfunctional technology (Liban, 2018; Vaithianathan et al., 2019; Ross, 2021) or creating better infrastructure or policies around its use (Dastin, 2020; Sisitzky and Schaefer, 2021; Ross, 2021). For example, when independent investigations revealed flaws in Epic’s sepsis prediction model (which was deployed in several hundred U.S. hospitals) (Wong et al., 2021), the company completely “overhauled” its original algorithm and released a second version that used completely different input and target variables, and recommended that the model be trained on a hospital’s own data rather than being applied off-the-shelf (Ross, 2021). However, repairs that improved the technology by making it more functional or less biased did not redress harms that resulted from the algorithms’ existence.
@@
Abandoned by the algorithm owner’s choice. In our analysis, the majority of systems that were abandoned were abandoned by choice, i.e., by a decision made by an organization that developed or used the algorithm. To name a few, Korean start-up ScatterLab disabled their Lee-Luda chatbot (Kwon and Yun, 2021), Twitter decided to stop using an algorithm to crop users’ images (Chowdhury, 2021), and leadership at Oregon’s Department of Human Services decided to stop using a child welfare call screening algorithm (ap2, 2022). Relatively few algorithm owners released public statements with explanations or named reasons why they decided to abandon the algorithm. Public statements that we found often referenced algorithmic harms as a primary reason why the system was abandoned (Lee, 2016; Chowdhury, 2021; Kwon and Yun, 2021; Hoover, 2023).
@@
Abandoned through policy. Algorithms overseen by uncooperative algorithm owners could still be abandoned through another governing body creating or enforcing applicable laws. Several algorithms in our case set were deemed illegal through the interpretation of existing laws (including laws written without an algorithm in mind) in a court of law. For example, a Dutch court ruled that their government must stop using an algorithm that falsely predicted that thousands of low-income residents had committed benefits fraud, after finding that the algorithm violated citizens’ right to privacy in the European Convention on Human Rights (aw2, 2020). U.S. courts ruled Medicaid benefits allocation algorithms in violation of existing laws (i.e., due process law and the Administrative Procedure Act) in both Arkansas and Idaho (Calo and Citron, 2021; Eppink, 2023). Other algorithms were banned by existing governing bodies founded to protect citizens’ rights, such as the U.S. Federal Trade Commission issuing a five-year ban on Rite Aid’s use of facial recognition because of the technology’s risks to consumers (ftc, 2023).
However, all past cases of systems abandoned through executive or judicial enforcement took a significantly longer time (e.g., several years (Liban, 2018; Eppink, 2023)) before the decision to abandon the algorithm was reached. These cases also required significant investment: e.g., the legal team representing Idaho Medicaid beneficiaries spent over $40,000 on expert consultations and 2,000 attorney hours to secure a settled agreement (Eppink, 2023).